# Discriminative Fine-tuning of LVLMs

### v2

@@ 3.4 How does the model’s behavior change?
Building upon the analysis from Sec. 3.1, we show that our training approach elicits the following behavioral changes: (1) The summary token - vision tokens’ attention increases in density. (2) Both the entropy of the output distribution of the summary token and the spread of the cumulative variance of the embeddings increase.
@@ 4 Experiments
Models compared: We compare with state-of-the-art models based on the two-towers (independent) approach, which is practical for retrieval purposes and also followed by our method. We cover a wide variety of settings, including different models and model sizes, training data, training losses, etc.: CLIP (ViT-B and ViT-L[36] -- the original CLIP trained with a contrastive loss on 400M image-text pairs; BLIP (ViT-L[28] -- trained on over 120M samples using contrastive, captioning and image-text matching losses; BLIP2 (T5-XXL) -- improved and scaled-up version of BLIP that makes use of qformer and LLM; OpenCLIP (ViT-G/14[37] -- scaled-up version of [36] trained on 2B samples; OpenCLIP (ViT-BigG/14[37], EVA-02-CLIP (ViT-E/14+[38], EVA-CLIP (8B[39] and EVA-CLIP (18B[39] -- large contrastively trained models, with up to 18B parameters, fine-tuned from vision encoders trained with Masked Image Modeling (MIM); E5-V (LLaVA-Next-8B) and E5-V (LLaVA-1.5-7B) -- LVLMs finetuned using a text-text contrastive loss. Depending on the task, we also include additional specialized baselines (e.g. NegCLIP [47] for compositionality).
@@ 4.1 Zero-shot image-text retrieval
Across all three datasets, our approach significantly surpasses the current state-of-the-art models, including models of similar size. It even outperforms the much bigger EVA-CLIP (18B) model (85.0% vs. 83.3%) on Flickr30k, (59.0% vs. 55.6%) on MS-COCO and (72.3% vs. 69.3%) on nocaps in terms of @R1 for image retrieval. Similarly, we also outperform the LVLM-based E5-V model by 5.5% on Flickr30k, 6% on MS-COCO, and 6.2% on nocaps.
@@ 4.2 Image-text compositionality
Compositionality benchmarks construct hard negatives that elicit various types of compositional failures. Herein, we focus our comparison on the currently most challenging test sets, SugarCrepe [19] and SugarCrepe++ [16] (For Winoground [41] please see supp. material). For SugarCrepe++, we are primarily interested in the Image-to-Text (ITT) setting since the Text-to-Text (TOT) setting evaluates the language component of the methods only.
As Tables 2 and 3 show, our approach is the best in both SugarCrepe and SugarCrepe++ (ITT). On SugarCrepe, we outperform the 18B EVA-CLIP model on all categories, with particularly large gains on relation replacement (76.1 vs. 86.8), attribution adding (85.0 vs. 95.8), and object swapping (65.3 vs. 85.8). The last case is particularly interesting as it directly measures the bag-of-words behavior, showcasing the significant improvements our method achieves in this direction. Additionally, we outperform the E5-V variant based on the same LLaVA-1.5-7B model that we used, and the one based on the heavier LLaVA-Next-8B. A similar trend is observed on SugarCreppe++ where we outperform EVA-CLIP (18B) by up to 10.9% (on object swapping) and E5-V (ITT) in all but relation replacement. Note that, thanks to its text-text training, E5-V surpasses our method for the TOT setting, but we note that their loss can be readily incorporated into our framework, leaving this for future work.
@@ 5.1 Impact of method’s components
We quantify the impact of the proposed method’s components by training on a smaller 1M subset, reporting results on SugarCrepe and on Flickr30k.
@@
While LoRA fine-tuning performs better than soft-prompting (due to its bigger capacity), the latter alone performs surprisingly well. To understand why, we analyze the changes the soft prompts undergo by finding the closest embedding in the LLM’s vocabulary. This results in the following decoded sentences: “</s> ’<Summarize the provided image in one word:/ \$[” and, “ $\omega$aSummarize the provided text in one word:--”. The two sentences remain unchanged semantically, with the only characters changed being the ones at the start and end of the prompt. Intuitively, this allows the model to mark/specialize the token that should gather the visual or textual evidence for discriminative tasks.
@@ 7 Results for additional model sizes and architectures
To further showcase the generalizability of our approach, herein we report results on two additional models: LLaVA-1.5-13B [34] and Qwen2-VL-2B [43]. The 1st is a scaled-up version of the LLaVA-1.5-7B [34] used in the main manuscript and tests the scalability of our approach with size. The second follows a different architecture and training procedure and has “only” 2B parameters, testing both generalizations to different architectures and finetuning in a lower-parameters regime. As the results from Tables 56 7 and 8 show, on all 6 datasets (i.e. Flickr, coco, nocaps, SugarCrepe, SugarCrepe++ and Winoground) for both retrieval and compositionality, in all cases we significantly improve upon the original zero-shot model performance, showing good scalability with size in both directions, i.e. for smaller and bigger models.
@@ 9 Zero-shot image recognition on ImageNet
Model Data. size Top-1 Top-5
CLIP (ViT-B[36] 400M 68.4 91.9
CLIP (ViT-L[36] 400M 74.0 94.0
EVA-CLIP (18B[39] 2.7B 83.5 97.2
CLIP (ViT-B[36] 15M 32.8 -
HiDeCLIP (ViT-B[36] 15M 45.9 -
FFF (ViT-B[7] 15M 51.1 -
BLIP (ViT-L[28] 129M 54.2 81.5
BLIP2 (ViT-L[29] 129M 46.7 74.2
LLaVA-Next-8B [26] 0M 45.8 74.6
E5-V [23] (LLaVA-Next-8B) 0M 48.2 76.6
LLaVA-1.5-7B [34] 0M 42.0 74.6
VladVA (Ours) (LLaVA-1.5-7B) 8.1M 63.7 88.3
Qwen2-VL-2B [43] 0M 54.7 79.4
VladVA (Ours) (Qwen2-VL-2B) 8.1M 70.6 91.1

### v3

@@ 3.4 How does the model’s behavior change?
Building upon the analysis from Sec. 3.1, we show that our training approach elicits the following behavioral changes: (1) The attention map between the summary and vision tokens increases in density. (2) Both the entropy of the output distribution of the summary token and the spread of the cumulative variance of the embeddings increase.
@@ 4 Experiments
Models compared: We compare with state-of-the-art models based on the two-towers (independent) approach, which is practical for retrieval purposes and also followed by our method. We cover a wide variety of settings: different models and model sizes, training data, training losses, etc.: CLIP (ViT-L[42] -- the original CLIP trained with a contrastive loss on 400M image-text pairs; BLIP (ViT-L[31] -- trained on over 120M samples using contrastive, captioning and image-text matching losses; BLIP2 (T5-XXL) -- improved and scaled-up version of BLIP; OpenCLIP (ViT-G/14[44] -- scaled-up version of [42] trained on 2B samples; OpenCLIP (ViT-BigG/14[44], EVA-02-CLIP (ViT-E/14+[46], EVA-CLIP (8B[47] and EVA-CLIP (18B[47] -- large contrastively trained models, with up to 18B parameters, fine-tuned from vision encoders trained with Masked Image Modeling (MIM); E5-V (LLaVA-Next-8B) and E5-V (LLaVA-1.5-7B) -- LVLMs finetuned using a text-text contrastive loss. Depending on the task, we also include additional specialized baselines (e.g. NegCLIP [55] for compositionality).
@@
Training details: We use a LLaVA-1.5 (7B) [38] model due to its popularity and simplicity (for other models, see supp. material). For LoRA adapters, we set the rank and $\alpha$ to 16. The number of soft prompts is aligned to the length of the tokenized hand-crafted prompt. Unless otherwise stated, we train the models for 7 epochs, using a batch size of 1024, a learning rate of $41e-$ 4, no weight decay, and AdamW [39] optimizer with default values for $\beta_{1}$ and $\beta_{2}$. During training, the learning rate is decayed according to a cosine scheduler [40]. Depending on the data configuration, we use up to 32 A100 GPUs. All our models and training procedures were implemented using PyTorch [41] and DeepSpeed [43].
We used the following training data: a 4M random subset of OpenImages [26], CC3M ($\sim$2.8M images) [45], and ShareGPT-4V [9]. As no captions are available for OpenImages, we automatically label them with 5 captions using BLIP2 [32]. During training, only one caption is sampled at a time. For longer captions, we directly use the ShareGPT-4V [9] data, which we extend with synthetic short captions produced by BLIP2 in order to enable the training procedure proposed in Sec. 3.2.3. Similarly, CC3M is automatically annotated with long captions using ShareGPT4-V [9].
@@ 4.1 Zero-shot image-text retrieval
As shown in Tab. 6, across all three datasets, our approach significantly surpasses the current state-of-the-art including models of similar size. It even outperforms the much bigger EVA-CLIP (18B) model (85.0% vs. 83.3%) on Flickr30k, (59.0% vs. 55.6%) on MS-COCO and (72.3% vs. 69.3%) on nocaps in terms of @R1 for image retrieval. Similarly, we outperform the LVLM-based E5-V model by 5.5% on Flickr30k, 7% on MS-COCO, and 6.4% on nocaps.
@@ 4.2 Image-text compositionality
Herein, we focus our comparison on the currently most challenging test sets, SugarCrepe [20] and SugarCrepe++ [17] (for Winoground [49] please see supp. material). For SugarCrepe++, we are mostly interested in the Image-to-Text (ITT) setting since the Text-to-Text (TOT) one evaluates the language component of the methods only.
As Tabs. 2 and 3 show, our approach is the best in both SugarCrepe and SugarCrepe++ (ITT). On SugarCrepe, we outperform the 18B EVA-CLIP model on all categories, with particularly large gains on relation replacement (76.1 vs. 86.8), attribution adding (85.0 vs. 95.8), and object swap (65.3 vs. 79.0). The last case is particularly interesting as it directly measures the bag-of-words behavior, showcasing significant improvements offered by our method. Additionally, we outperform the E5-V variant based on the same LLaVA-1.5-7B model that we used, and the one based on the heavier LLaVA-Next-8B. A similar trend is observed on SugarCreppe++ where we outperform EVA-CLIP (18B) by up to 10.9% (on object swap) and E5-V (ITT) in all but relation replacement. Thanks to its text-text training, E5-V surpasses our method for the TOT setting, but we note that their loss can be readily incorporated into our framework, leaving this for future work.
@@ 5.1 Impact of method’s components
We quantify the impact of the proposed method’s components by training on a smaller 1M subset, reporting results on SugarCrepe (averaged over each category) and on Flickr30k (R@1 for T2I and I2T).
@@
While LoRA fine-tuning performs better than soft-prompting (due to its bigger capacity), the latter alone performs surprisingly well. To understand why, we analyze the changes the soft prompts undergo by finding the closest embedding in the LLM’s vocabulary. This results in the following decoded sentences: “</s> ’<Summarize the provided image in one word:/ \$[” and, “ $\omega$aSummarize the provided text in one word:--”. The two sentences remain unchanged semantically, with the only characters changed being the ones at the start and the end of the prompt. Intuitively, this allows the model to mark/specialize the token that should gather the visual or textual evidence for discriminative tasks.
@@
Table 5: Impact of training data size.
Training data SugarCrepe Flickr30k
Replace Swap Add T2I I2T
LLava-1.5-7B (0M) 81.9 59.8 64.7 59.6 65.6
OpenImages (1M) 87.1 72.0 88.6 79.6 92.9
OpenImages (4M) 88.2 79.6 89.1 82.3 93.1
$+$ ShareGPT-4V (1.3M) 91.0 80.3 92.4 83.1 94.0
        $+$ CC3M (2.8M) 92.3 80.9 95.5 85.0 94.3
5.2 Impact of training dataset size
Although at a relatively small scale (training is expensive due to the LVLM), herein, we aim to examine whether scaling the dataset size benefits the proposed discriminative adaptation of LVLMs. Specifically, we scale our dataset size from 1M to 8.1M samples. As Tab. 5 shows, we obtain steady gains across all metrics, with no signs of immediate saturation. This suggests that some potential is still left untapped, and further scaling could result in extra gains.
@@ 7 Results for additional model sizes and architectures
To further showcase the generalizability of our approach, herein we report results on two additional models: LLaVA-1.5-13B [37] and Qwen2-VL-2B [51]. The 1st is a scaled-up version of the LLaVA-1.5-7B [37] used in the main manuscript and tests the scalability of our approach with size. The second follows a different architecture and training procedure and has “only” 2B parameters, testing both generalizations to different architectures and finetuning in a lower-parameters regime. As the results from Tables 67 8 and 9 show, on all 6 datasets (i.e. Flickr, coco, nocaps, SugarCrepe, SugarCrepe++ and Winoground) for both retrieval and compositionality, in all cases we significantly improve upon the original zero-shot model performance, showing good scalability with size in both directions, i.e. for smaller and bigger models.
@@ 9 Zero-shot image recognition on ImageNet
Model Data. size Top-1 Top-5 CLIP (ViT-B[42] 400M 68.4 91.9 CLIP (ViT-L[42] 400M 74.0 94.0 EVA-CLIP (18B[47] 2.7B 83.5 97.2 CLIP (ViT-B[42] 15M 32.8 - HiDeCLIP (ViT-B[42] 15M 45.9 - FFF (ViT-B[8] 15M 51.1 - BLIP (ViT-L[31] 129M 54.2 81.5 BLIP2 (ViT-L[32] 129M 46.7 74.2 LLaVA-Next-8B [29] 0M 45.8 74.6 E5-V [24] (LLaVA-Next-8B) 0M 48.2 76.6 LLaVA-1.5-7B [37] 0M 42.0 74.6 VladVA (Ours) (LLaVA-1.5-7B) 8.1M 63.7 88.3 Qwen2-VL-2B [51] 0M 54.7 79.4 VladVA (Ours) (Qwen2-VL-2B) 8.1M 70.6 91.1
@@
10 Which layer to choose the token from?
In the main paper, we’ve used the last token of the last layer as the summary, discriminative token. Intuitively, by selecting the last layer, we maximize the amount of parameters we can adapt, and hence adaptation plasticity. However, herein, for completeness, we report results for different layer IDs in Table 11. The results show that the last 3-4 layers have comparable performance, performance that degrades as we select earlier layers.
Table 11: Performance change when using different layer IDs, reported on SugarCrepe (averaged) and Flickr30k (I2T).
Dataset/Layer 32 (last) 31 28 24 20 16
Flickr30K 0 +0.3 -0.2 -1.1 -13.5 -59.0
SugarCrepe: 0 +0.9 +0.4 -0.1 -8.1 -20.0